Biochemical Corp. requires $600,000 in financing over the next three years. The firm can borrow the funds for three years at 10.80 percent interest per year. The CEO decides to do a forecast and predicts that if she utilizes short-term financing instead, she will pay 7.50 percent interest in the first year, 12.15 percent interest in the second year, and 8.25 percent interest in the third year. Assume interest is paid in full at the end of each year.

a. Determine the total interest cost under each plan.
Long term fixed rate:

Short term fixed rate:



b. Which plan is less costly?
  

Long-term fixed-rate plan
Short-term variable-rate plan
Answer #1

Long term fixed rate plan:

Interest cost = 600,000 * 10.80% * 3

Interest cost = 194,400

Short term fixed rate plan:

Interest cost 1st year = 600,000 * 7.50%

Interest cost 1st year = 45,000

Interest cost 2nd year = 600,000 * 12.15%

Interest cost 2nd year = 72,900

Interest cost 3rd year = 600,000 * 8.25%

Interest cost 3rd year = 49,500

Total Interest Cost = 45,000 + 72,900 + 49,500

Total Interest Cost = 167,400

Part B

Short term fixed cost plan has less costly interest rates